SPECTROSCOPIC STRUCTURAL INVESTIGATIONS OF CHARGE-TRANSFER COMPLEXES OF n-DONORS AND SIGMA- AND PI-ACCEPTORS

\underline{EL-METWALLY NOUR}, SIHAM Y. ALQARADAWI, Department of Chemistry and Earth Sciences , Qatar University, P.O.Box 2713 Doha, Qatar.

Molecular charge transfer complexes of n-electron donors such as amines, pyridines and polynitrogen cyclic bases with both sigma- and pi-acceptors have been studied spectrophotometrically in chloroform. The reaction stoichiometries were determined using photometric titration methods. The formed CT-complexes are characterized and formulated based on electronic and infrared spectra as well as elemental and thermal measurements.